40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 10 22:17:29 2014, Koji, Summary, LSC, X/Y green beat mode overlap measurement 
    Reply  Mon Jun 22 12:57:09 2015, ericq, Summary, LSC, X/Y green beat mode overlap measurement redone 
       Reply  Mon Jun 22 14:53:37 2015, rana, Summary, LSC, X/Y green beat mode overlap measurement redone 
          Reply  Fri Oct 28 15:46:29 2016, gautam, Summary, LSC, X/Y green beat mode overlap measurement redone 
Message ID: 12587     Entry time: Fri Oct 28 15:46:29 2016     In reply to: 11370
Author: gautam 
Type: Summary 
Category: LSC 
Subject: X/Y green beat mode overlap measurement redone 

I've been meaning to do this analysis ever since putting in the new laser at the X-end, and finally got down to getting all the required measurements. Here is a summary of my results, in the style of the preceeding elogs in this thread. I dither aligned the arms and maximized the green transmission DC levels, and also the alignment on the PSL table to maximize the beat note amplitude (both near and far field alignment was done), before taking these measurements. I measured the beat amplitude in a few ways, and have reported all of them below...

             XARM   YARM 
o BBPD DC output (mV), all measured with Fluke DMM
 V_DARK:     +1.0    +3.0
 V_PSL:      +8.0    +14.0
 V_ARM:      +175.0  +11.0


o BBPD DC photocurrent (uA)
I_DC = V_DC / R_DC ... R_DC: DC transimpedance (2kOhm)

 I_PSL:       3.5    5.5
 I_ARM:      87.0    4.0


o Expected beat note amplitude
I_beat_full = I1 + I2 + 2 sqrt(e I1 I2) cos(w t) ... e: mode overlap (in power)

I_beat_RF = 2 sqrt(e I1 I2)

V_RF = 2 R sqrt(e I1 I2) ... R: RF transimpedance (2kOhm)

P_RF = V_RF^2/2/50 [Watt]
     = 10 log10(V_RF^2/2/50*1000) [dBm]

     = 10 log10(e I1 I2) + 82.0412 [dBm]
     = 10 log10(e) +10 log10(I1 I2) + 82.0412 [dBm]

for e=1, the expected RF power at the PDs [dBm]
 P_RF:      -13.1  -24.5


o Measured beat note power (measured with oscilloscope, 50 ohm input impedance)      
 P_RF:      -17.8dBm (81.4mVpp)  -29.8dBm (20.5mVpp)   (38.3MHz and 34.4MHz)  
    e:        34                    30  [%]                          
o Measured beat note power (measured with Agilent RF spectrum analyzer)       
 P_RF:      -19.2  -33.5  [dBm] (33.2MHz and 40.9MHz)  
    e:       25     13    [%]                          

I also measured the various green powers with the Ophir power meter: 

o Green light power (uW) [measured just before PD, does not consider reflection off the PD]
 P_PSL:       16.3    27.2
 P_ARM:       380     19.1

Measured beat note power at the RF analyzer in the control room
 P_CR:      -36    -40.5    [dBm] (at the time of measurement with oscilloscope)
Expected    -17    - 9    [dBm] (TO BE UPDATED)

Expected Power: (TO BE UPDATED)
Pin + External Amp Gain (25dB for X, Y from ZHL-3A-S)
    - Isolation trans (1dB)
    + GAV81 amp (10dB)
    - Coupler (10.5dB)


The expected numbers for the control room analyzer in red have to be updated. 

The main difference seems to be that the PSL power on the Y broadband PD has gone down by about 50% from what it used to be. In either measurement, it looks like the mode matching is only 25-30%, which is pretty abysmal. I will investigate the situation further - I have been wanting to fiddle around with the PSL green path in any case so as to facilitate having an IR beat even when the PSL green shutter is closed, I will try and optimize the mode matching as well... I should point out that at this point, the poor mode-matching on the PSL table isn't limiting the ALS noise performance as we are able to lock reliably...

ELOG V3.1.3-